OpenStudy (anonymous):

what is the max area of an isosceles triangle with two side lengths of 5? Use calculus optimization of max and min please.

OpenStudy (anonymous):

\[\sqrt{3}\times6.25\]

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (anonymous):

u know the fact area of an eqilateral trngle s most

OpenStudy (anonymous):

its not equilateral though?

OpenStudy (anonymous):

or u can use max. or min. concept

OpenStudy (anonymous):

y not

OpenStudy (anonymous):

yeah, but how do you use the max or min concept??

OpenStudy (anonymous):

its says its an isosceles.

OpenStudy (anonymous):

take the base x and figure rhe area with the help of x then difrntiate it

OpenStudy (anonymous):

eqilateral is also isosceles

OpenStudy (anonymous):

but is the equilateral area the MAXIMUM area it could be? thats the thing.. and i can take the base and find the area and differentiate..but i can't get it to come out to an answer...

OpenStudy (cwrw238):

u can use area = 0.5 *5*5 * sin x where x = vertex angle (if i remember the formula crrectly)

OpenStudy (anonymous):

okay this is what i have area= .5(x)((25-((x^2)/4))^(1/2)) when i try to differntiate that is when i get stuck

OpenStudy (anonymous):

use pythagorus

OpenStudy (anonymous):

i dont know its a right triangle! i cant use pythag.

OpenStudy (anonymous):

den do diffrntiate and see it comes that

OpenStudy (anonymous):

the diffientiation is what im getting stuck at.

OpenStudy (anonymous):

join the vertex to mid pnt of base then apply pyth.

OpenStudy (anonymous):

okayy. that is not getting me the max. i have to use CALCULUS.

OpenStudy (anonymous):

who restrict u

OpenStudy (anonymous):

my teacher..

OpenStudy (cwrw238):

A = (25-x^2)^(1/2) * x where x = half base dA/dx = (25-x^2)^(1/2) + x* (1/2)* (25-x^2)^(-1/2)* -2x [by product rule) dA/dx = (25-x^2)^(1/2) - x^2 / (25-x^2)^(1/2) = 25 - x^2 - x^2 = 0 for max/min x = sqrt 12.5 max area is when length of base = 2* sqrt 12.5 max area = height * sqrt 12.5 = (25-12.5)^(1/2) * sqrt 12.5